Sandro Clean uses a job order cost accounting system. On October 1, the company has a balance in Work in Process Inventory of $4,200 and two jobs in process: Job No. R92, $1,600 and Job No. R93, $2,600. During October, a summary of source documents reveals the following:

Job Number Materials Requisitions Slips Labor Time Ticket
Job No. R92 $2,200

$7,1000

Job No. R93 1,700 4,100
Job No. R94 4,700 3,300
Job No.R95 2,200 $5,100
General Use 1,800 2,000
Total $12,600 $21,600

Sandro applies manufacturing overhead to jobs at an overhead rate of 90% of direct labor cost. Job No. R92 was completed during the month.

Instructions

(a) Prepare summary journal entries to record the requisition slips, time tickets, the assignment of manufacturing overhead to jobs, and the completion of Job No. R92. Show computations.

(b)    Answer the following questions.

   1.   What is the balance in Work in Process Inventory at October 31?

   2.   If Sandro incurred $13,000 of manufacturing overhead in addition to indirect labor and indirect materials, was overhead over- or underapplied in October and by how much?
Answer #1

a) Journal entry

No General Journal Debit Credit
1 Work in process 10800
Manufacturing overhead 1800
Raw material 12600
2 Work in process 19600
Manufacturing overhead 2000
Wages payable 21600
3 Work in process (19600*90%) 17640
Manufacturing overhead 17640
4 Finished goods 17290
work in process 17290

b1) Ending work in process = 8600+12500*1.9+2600 = 34950

b2) Applied overhead = 17640

Actual overhead = 13000+1800+2000 = 16800

Over applied overhead = 17640-16800 = 840